OpenStudy (laddiusmaximus):

Determine the values of x for which the function f is decreasing, where f(x)=(440−x)^(2)+x^(2) enter your answer in interval notation use your answer to determine which is bigger, 154^(2) + 286^(2) or 198^(2) + 242^(2)

OpenStudy (anonymous):

take the derivative to get f'(x) = 2(440 - x)(-1) + 2x = -880 + 4x. Find the zeros of the derivative to get x = 220. your intervals are (-inf, 220), (220, inf). pick a value in these intervals and plug in to the derivative, if negative, f is decreasing. if positive, f is increasing.

OpenStudy (laddiusmaximus):

find the zeros?

OpenStudy (precal):

solve for 0 ie -880+4x=0

OpenStudy (precal):

Actually it was done for you 220

OpenStudy (precal):



OpenStudy (precal):

Pick a number from the intervals, and test it to see if it is positive or negative

OpenStudy (laddiusmaximus):

i plug the number into 4x-880?

OpenStudy (precal):